A Bibliography of NATO Advanced Study Institute Proceedings

Karl Berry 135 Center Hill Rd. Plymouth, MA 02360

USA

E-mail: karl@cs.umb.edu (Internet)

25 October 2010 Version 1.35

Abstract

This bibliography records publications in the NATO Advanced Study Institute Proceedings.

Title word cross-reference

L^{1} [Sha89]. n [Bjö89]. R^{n} [Gab91, Eva89]. Z_{n} [Bjö89].

-roots [Bjö89].

Absolutely [Art89]. accurate [SA91]. Acoustics [Sch89b]. Advanced [BB90]. Advances [BB90]. analyses [Mad91]. Analysis [Bro89]. Analytic [Clu89]. Anatomic [KFC89]. Antenna [Bro89]. Application [Klo89]. Applications [KL89]. Applied [Üns89]. Approach [Grö89]. Approximations [Sei89]. Architectures [Sch89a]. Atomic [Hay89]. attractors [CM91]. Auto [Tüz89]. Auto-Cross [Tüz89].

Band [Grö89]. Band-Limited [Grö89]. bandlimited [Grö91]. Barrier [Cul89]. basic [Ble91]. behaviour [Saf91]. Bispectra [Pap89]. broadband [SA91].

Carlo [Kör91]. Cauchy [SA91]. Certain [Ben89b]. characteristic [Çap91]. Characterization [MH91]. Circle [NG89]. Class [Ben89b]. Classical [Ing89]. Coding [Sch89a]. Coefficients [NG89]. Coherent [Fei89]. Combinatorial [Ble89]. Combined [dHdV89]. Comparison [Gür89]. Complex [Üns89]. Computations [TA89]. Computer [Sch89b]. Conjugate [PSA89]. Constant [Bjö89]. continuations [Kah91]. control [CP91]. Convergent [Art89]. Correlation [Tüz89]. Cross [Tüz89]. Crystallography [Hau89]. Cyclic [Bjö89].

Data [Jak89]. decomposition [SKR^{+}91]. Decompositions [Hay89]. definite [Gab91]. Density [CM91]. Derivatives [Clu89]. differential [Jaf91]. dilation [Hei91]. Dimension [Ble89]. dimensions [CP91]. Diodes [Cul89]. Diophantine [Dod89].

Direct [Hau89]. Distributions [Art89]. divisibility [Car91]. Domain [Tij89]. dynamical [Bro91].

Effects [Cul89]. Electromagnetic [Tij89]. Electromagnetics [PSA89]. Empirical [Çap91]. Energy [Üns89]. entropy [BK91]. environmental [Chr91]. equations [CP91]. equivalent [Kou91]. Estimate [Sha89]. Estimation [Ben89a]. Euclidean [KL89]. Even [Tüz89]. Expansions [HW89]. Extension [Gab91]. Extensions [CLH89]. Extracting [HSH89].

factorization [Pap91]. FFT [PSA89]. fields [CL91]. Finite [Üns89]. Fixed [KL91]. Fourier [Bjö89]. fractals [CM91]. Frames [Fei89]. Function [HSH89]. Functional [Gür89]. Functions [Art89]. fuzzy [CM91].

Gabor [HW89, SKR^{+}91]. Generalized [HSH89]. Generation [SA91]. geometries [Kee91]. geometry [PR91]. Gradient [PSA89]. Groupoids [Sei89]. Groups [Kar89]. Growth [Jak89].

Hamilton [CP91]. Hamilton-Jacobi [CP91]. harmonic [Cul89]. High [Lah89]. High-resolution [Lah89]. Holographic [Sch89a]. holography [Sch91].

Identification [Pap89]. Image [Sch89a]. images [SKR^{+}91]. Inequalities [Ben89a]. inference [Çap91]. infinite [CP91]. information [SA91]. Innovations [Pap91]. Institute [BB90]. integration [Ble91]. Interdependencies [Ble89]. Interpolation [Jak89]. Invariance [Roy89]. Invariant [Mad89]. inverse [CM91]. Irregular [Fei89]. iterated [PR91].

Jacobi [CP91].

Laplace [dHdV89]. Layered [dHdV89]. Limit [Ing89]. Limited [Grö89]. Linear [SKR^{+}91].

mathematical [New91]. Maximum [Ben89b]. Measurements [Ble89]. measures [BK91]. Media [dHdV89]. Method [HSH89]. Methods [Hau89]. Metrizable [Kar89]. Minimal [Hau89]. mixed [KL91]. Modulus [Ben89b]. Monte [Kör91]. Multiband [Dod89]. Multifractal [Pey91]. Multiresolution [Mad91]. Multiscale [Mad89].

narrowband [SA91]. NATO [BB90]. networks [Nan91]. neural [Nan91]. Neurocomputer [Sch89a]. non [KL89]. non-Euclidean [KL89]. non-linear [SKR^{+}91]. Normal [Bro91]. Nuclear [Kal91]. Number [Sch89b]. numbers [Bro91]. Numerical [Tij89].

Odd [Tüz89]. Operator [Sha89]. optimal [CP91]. order [CP91]. organising [Nan91]. other [CM91].

Paley [Kah91]. paper [Nan91]. part [New91]. partial [Jaf91]. Parts [Tüz89]. pattern [Nan91]. Pencil [HSH89]. Pencil-of-Function [HSH89]. periodograms [Kör91]. phase [Saf91]. Physics [KL89]. Plancherel [Eva89]. points [KL91]. Poles [HSH89]. Polynomials [Ben89b]. positive [Gab91]. positive-definite [Gab91]. Power [Jak89]. prediction [Pap91]. Principle [Ben89a]. probabilistic [Nan91]. probability [New91]. problem [CM91]. Problems [PSA89]. Proceedings [BB90]. Processing [PSA89]. Products [Kar89]. Propagation [dHdV89]. Properties [NG89].

Quantum [Sch91].

radial [Gab91]. Radon [Wal91]. random [CL91]. rate [Pap91]. Rates [Dod89]. ray [Hau89]. recognition [Nan91]. Recurrence [Lah89]. Reflector [Bro89]. Relations [Lah89]. research [Chr91]. resolution [Lah89]. Responses [HSH89]. Restriction [CLH89]. Results [Art89]. Riesz [Kar89]. roots [Bjö89].

Sampling [Dod89]. scaling [Mad91]. Scattering [Tij89]. Schottky [Cul89]. Schottky-Barrier [Cul89]. Science [Sch89b]. Second [Cul89]. Second-harmonic [Cul89]. self [Nan91]. self-organising [Nan91]. Semi [Ing89]. Semi-Classical [Ing89]. sequence [Çap91]. Series [Art89]. sets [CM91]. Shape [KFC89]. Sharp [Grö91]. Signal [PSA89]. Signals [Dod89]. singular [Car91]. singularities [MH91]. Solution [Tij89]. solving [Hei91]. some [Gür89]. Space [KL89]. space-valued [Kal91]. Spaces [Gür89]. spatiotemporal [Chr91]. Spectral [Lah89]. Spectras [Gür89]. Spectrum [Ben89a]. Spline [Jak89]. Stationary [Ben91]. step [CM91]. stochastic [BK91]. Storage [Üns89]. Structures [KFC89]. Study [BB90]. Successive [Clu89]. surface [Kee91]. Synthesis [Bro89]. System [Pap89]. systems [Bro91].

Techniques [Roy89]. their [Chr91]. Theorem [Eva89]. Theorems [CLH89]. Theory [dHdV89]. Thermal [Üns89]. tiles [Mad91]. Time [Tij89]. Time-Domain [Tij89]. Transform [CLH89]. Transformation [Klo89]. Transformations [dHdV89]. Transforms [Bjö89]. Transient [dHdV89].

Translation [Mad89].

Trigonometric [Ben89b]. truth [New91]. tutorial [Nan91].

ultraflat [Saf91]. Uncertainty [Ben89a]. Unimodular [Ben89b]. Unit [NG89]. Using [Lah89].

valued [Kal91]. view [Ble91].

Wave [dHdV89]. waveforms [Car91]. Wavelet [HW89]. Wavelets [Jaf91]. Weighted [CLH89]. Whose [Bjö89]. Wiener [Eva89]. Wiener-Plancherel [Eva89]. work [Kah91].

X [Hau89]. X-ray [Hau89].

Zygmund [Kah91].

References

Artemiadis:FAA-89-311

[Art89] N. K. Artémiadis. Results on the absolutely convergent series of functions and of distributions. In Byrnes and Byrnes [BB90], pages 311–316. ISBN 0-7923-0875-1.

Byrnes:FAA-89

[BB90] J. S. Byrnes and Jennifer L. Byrnes, editors. Recent Advances in Fourier Analysis and its Applications: Proceedings of the NATO Advanced Study Institute, volume 315 of NATO ASI series C: Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0875-1.

Byrnes:PSM-91

[BBBH92] J. S. Byrnes, Jennifer L. Byrnes, Karl Berry, and Kathryn A. Hargreaves, editors. Probabilistic and Stochastic Methods in Analysis, with Applications: Proceedings of the NATO Advanced Study Institute, volume 372 of NATO ASI series C: Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1804-8.

Benedetto:FAA-89-143

[Ben89a] John J. Benedetto. Uncertainty principle inequalities and spectrum estimation. In Byrnes and Byrnes [BB90], pages 143–182. ISBN 0-7923-0875-1.

Benke:FAA-89-83

[Ben89b] George Benke. On the maximum modulus for a certain class of unimodular trigonometric polynomials. In Byrnes and Byrnes [BB90], pages 83–100. ISBN 0-7923-0875-1.

Benedetto:PSM-91-117

[Ben91] John J. Benedetto. Stationary frames and spectral estimation. In Byrnes et al. [BBBH92], pages 117–163. ISBN 0-7923-1804-8.

Bjorck:FAA-89-131

[Bjö89] Göran Björck. Functions of modulus 1 on Z_{n} whose Fourier transforms have constant modulus, and “cyclic n-roots”. In Byrnes and Byrnes [BB90], pages 131–140. ISBN 0-7923-0875-1.

Bisbas:PSM-91-663

[BK91] A. Bisbas and C. Karanikas. On the entropy of certain stochastic measures. In Byrnes et al. [BBBH92], pages 663–668. ISBN 0-7923-1804-8.

Blei:FAA-89-631

[Ble89] Ron C. Blei. Combinatorial dimension and measurements of interdependencies. In Byrnes and Byrnes [BB90], pages 631–642. ISBN 0-7923-0875-1.

Blei:PSM-91-649

[Ble91] Ron C. Blei. A basic view of stochastic integration. In Byrnes et al. [BBBH92], pages 649–662. ISBN 0-7923-1804-8.

Brown:FAA-89-367

[Bro89] Roy C. Brown. Fourier analysis in reflector antenna synthesis. In Byrnes and Byrnes [BB90], pages 367–376. ISBN 0-7923-0875-1.

Brown:PSM-91-207

[Bro91] Gavin Brown. Normal numbers and dynamical systems. In Byrnes et al. [BBBH92], pages 207–216. ISBN 0-7923-1804-8.

Capar:PSM-91-517

[Çap91] Uluğ Çapar. Empirical characteristic functional analysis and inference in sequence spaces. In Byrnes et al. [BBBH92], pages 517–534. ISBN 0-7923-1804-8.

Carasso:PSM-91-273

[Car91] Alfred S. Carasso. Infinite divisibility and the identification of singular waveforms. In Byrnes et al.

[BBBH92], pages 273–286. ISBN 0-7923-1804-8.

Christakos:PSM-91-287

[Chr91] George Christakos. Certain results on spatiotemporal random fields and their applications in environmental research. In Byrnes et al. [BBBH92], pages 287–322. ISBN 0-7923-1804-8.

Chou:PSM-91-343

[CL91] I-Shang Chou and Keh-Shin Lii.

Spectral analysis of random fields with random sampling. In Byrnes et al. [BBBH92], pages 343–368. ISBN 0-7923-1804-8.

Carton-Lebrun:FAA-89-579

[CLH89] C. Carton-Lebrun and H. P.

Heinig. Weighted extensions of restriction theorems for the Fourier transform. In Byrnes and Byrnes [BB90], pages 579–596. ISBN 0-7923-0875-1.

Clunie:FAA-89-623

[Clu89] J. G. Clunie. Successive derivatives of analytic functions. In Byrnes and Byrnes [BB90], pages 623–630. ISBN 0-7923-0875-1.

Cabrelli:PSM-91-163

[CM91] Carlos Cabrelli and Ursula Molter.

Density of fuzzy attractors: A step towards the solution of the inverse problem for fractals and other sets. In Byrnes et al. [BBBH92], pages 163–173. ISBN 0-7923-1804-8.

Cannarsa:PSM-91-617

[CP91] P. Cannarsa and G. Da Prato. Second order Hamilton-Jacobi equations in infinite dimensions and stochastic optimal control problems. In Byrnes et al. [BBBH92], pages 617–630. ISBN 0-7923-1804-8.

Cullen:FAA-89-499

[Cul89] A. L. Cullen. Second-harmonic effects in Schottky-Barrier diodes. In Byrnes and Byrnes [BB90], pages 499–506. ISBN 0-7923-0875-1.

Hoop:FAA-89-387

[dHdV89] Adrianus T. de Hoop and Sytze M. de Vries. Combined Laplace and Fourier transformations in the theory of transient wave propagation in layered media. In Byrnes and Byrnes [BB90], pages 387–426. ISBN 0-7923-0875-1.

Dodson:FAA-89-483

[Dod89] M. M. Dodson. Diophantine inequalities and sampling rates for multiband signals. In Byrnes and Byrnes [BB90], pages 483–498. ISBN 0-7923-0875-1.

Evans:FAA-89-613

[Eva89] Ward R. Evans. The Wiener-Plancherel theorem on R^{n}. In Byrnes and Byrnes [BB90], pages 613–622. ISBN 0-7923-0875-1.

Feichtinger:FAA-89-427

[Fei89] Hans G. Feichtinger. Coherent frames and irregular sampling. In Byrnes and Byrnes [BB90], pages 427–440. ISBN 0-7923-0875-1.

Gabardo:PSM-91-543

[Gab91] Jean-Pierre Gabardo. Extension of radial positive-definite distributions in R^{n} and maximum entropy. In Byrnes et al. [BBBH92], pages 543–554. ISBN 0-7923-1804-8.

Grochenig:FAA-89-251

[Grö89] Karlheinz Gröchenig. A new approach to irregular sampling of band-limited functions. In Byrnes and Byrnes [BB90], pages 251–260. ISBN 0-7923-0875-1.

Grochenig:PSM-91-323

[Grö91] Karlheinz Gröchenig. Sharp results on irregular sampling of bandlimited functions. In Byrnes et al. [BBBH92], pages 323–336. ISBN 0-7923-1804-8.

Gurkanli:FAA-89-643

[Gür89] A. Turan Gürkanli. Comparison of spectras in some functional spaces. In Byrnes and Byrnes [BB90], pages 643–652. ISBN 0-7923-0875-1.

Hauptman:FAA-89-3

[Hau89] Herbert Hauptman. A minimal principle in the direct methods of X-ray crystallography. In Byrnes and Byrnes [BB90], pages 3–16. ISBN 0-7923-0875-1.

Hayman:FAA-89-597

[Hay89] W. K. Hayman. Atomic decompositions. In Byrnes and Byrnes [BB90], pages 597–612. ISBN 0-7923-0875-1.

Heil:PSM-91-15

[Hei91] Christopher Heil. Methods of solving dilation equations. In Byrnes et al. [BBBH92], pages 15–46. ISBN 0-7923-1804-8.

Heller:PSM-91-101

[Hel91] William Heller. Complex analysis and frames in sampling theory.

In Byrnes et al. [BBBH92], pages 101–116. ISBN 0-7923-1804-8.

Houdre:PSM-91-337

[Hou91] Christian Houdré. Some recent results on the sampling theorem. In Byrnes et al. [BBBH92], pages 337–342. ISBN 0-7923-1804-8.

Hua:FAA-89-183

[HSH89] Yingbo Hua, Tapan K. Sarkar, and Fengdo Hu. The generalized pencil-of-function method for extracting poles from transient responses. In Byrnes and Byrnes [BB90], pages 183–200. ISBN 0-7923-0875-1.

Heil:FAA-89-441

[HW89] Christopher Heil and David Walnut. Gabor and wavelet expansions. In Byrnes and Byrnes [BB90], pages 441–454. ISBN 0-7923-0875-1.

Ingolfsson:FAA-89-29

[Ing89] Ketill Ingolfsson. Harmonic analysis of the semi-classical limit. In Byrnes and Byrnes [BB90], pages 29–40. ISBN 0-7923-0875-1.

Jaffard:PSM-91-3

[Jaf91] Stéphane Jaffard. Wavelets and analysis of partial differential equations. In Byrnes et al. [BBBH92], pages 3–15. ISBN 0-7923-1804-8.

Jakimovski:FAA-89-73

[Jak89] A. Jakimovski. Spline interpolation of data of power growth. In Byrnes and Byrnes [BB90], pages 73–75. ISBN 0-7923-0875-1.

Kahane:PSM-91-481

[Kah91] Jean-Pierre Kahane. Some continuations of the work of Paley and Zygmund on random trigonometric series. In Byrnes et al. [BBBH92], pages 481–516. ISBN 0-7923-1804-8.

Kallianpur:PSM-91-631

[Kal91] Gopinath Kallianpur. Nuclear space-valued stochastic differential equations with applications. In Byrnes et al. [BBBH92], pages 631–648. ISBN 0-7923-1804-8.

Karanikas:FAA-89-653

[Kar89] C. Karanikas. On a class of Riesz products on metrizable groups. In Byrnes and Byrnes [BB90], pages 653–660. ISBN 0-7923-0875-1.

Keenan:PSM-91-469

[Kee91] Daniel Keenan. Random surface geometries with applications to image processing. In Byrnes et al. [BBBH92], pages 469–478. ISBN 0-7923-1804-8.

Kennedy:FAA-89-17

[KFC89] D. N. Kennedy, P. A. Filipek, and V. S. Caviness. Fourier shape analysis of anatomic structures. In Byrnes and Byrnes [BB90], pages 17–28. ISBN 0-7923-0875-1.

Kovalyov:FAA-89-41

[KL89] M. Kovalyov and M. Legare.

Fourier transform in non-Euclid-ean space and its applications to physics. In Byrnes and Byrnes [BB90], pages 41–58. ISBN 0-7923-0875-1.

Kedem:PSM-91-573

[KL91] Benjamin Kedem and Silvia Lopes. Fixed points in mixed spectrum analysis. In Byrnes et al. [BBBH92], pages 573–592. ISBN 0-7923-1804-8.

Klooster:FAA-89-221

[Klo89] K. Van’t Klooster. On the application of Fourier transformation in antenna measurements. In Byrnes and Byrnes [BB90], pages 221–236. ISBN 0-7923-0875-1.

Korner:PSM-91-593

[Kör91] T. W. Körner. Monte carlo periodograms. In Byrnes et al. [BBBH92], pages 593–604. ISBN 0-7923-1804-8.

Koumandos:PSM-91-605

[Kou91] Stamatis Koumandos. A class of equivalent measures. In Byrnes et al. [BBBH92], pages 605–614. ISBN 0-7923-1804-8.

Lahr:FAA-89-463

[Lah89] J. H. G. Lahr. High-resolution spectral analysis using recurrence relations. In Byrnes and Byrnes [BB90], pages 463–482. ISBN 0-7923-0875-1.

Madych:FAA-89-455

[Mad89] W. R. Madych. Translation invariant multiscale analysis. In Byrnes and Byrnes [BB90], pages 455–462. ISBN 0-7923-0875-1.

Madych:PSM-91-233

[Mad91] W. R. Madych. Multiresolution analyses, tiles, and scaling functions. In Byrnes et al. [BBBH92], pages 233–244. ISBN 0-7923-1804-8.

Mallat:PSM-91-47

[MH91] Stéphane Mallat and Wen Liang Hwang. Characterization of singularities. In Byrnes et al. [BBBH92], pages 47–100. ISBN 0-7923-1804-8.

Nandagopal:PSM-91-369

[Nan91] D. Nandagopal. Application of probabilistic and self-organising neural networks in pattern recognition: a tutorial paper. In Byrnes et al. [BBBH92], pages 369–382. ISBN 0-7923-1804-8.

Newman:PSM-91-535

[New91] Donald J. Newman. Is probability a part of mathematical truth? In Byrnes et al. [BBBH92], pages 535–542. ISBN 0-7923-1804-8.

Newman:FAA-89-79

[NG89] Donald J. Newman and André Giroux. Properties on the unit circle of polynomials with unimodular coefficients. In Byrnes and Byrnes [BB90], pages 79–82. ISBN 0-7923-0875-1.

Papoulis:FAA-89-377

[Pap89] Athanasios Papoulis. Bispectra and system identification. In Byrnes and Byrnes [BB90], pages 377–386. ISBN 0-7923-0875-1.

Papoulis:PSM-91-247

[Pap91] Athanasios Papoulis. Innovations and entropy rate with applications in factorization, spectral estimation, and prediction. In Byrnes et al. [BBBH92], pages 247–262. ISBN 0-7923-1804-8.

Peyriere:PSM-91-175

[Pey91] Jacques Peyrière. Multifractal measures. In Byrnes et al. [BBBH92], pages 175–186. ISBN 0-7923-1804-8.

Piccioni:PSM-91-669

[PR91] Mauro Piccioni and Massimo Regoli. The geometry of attractors for a class of iterated function systems. In Byrnes et al. [BBBH92], pages 669–686. ISBN 0-7923-1804-8.

Ponnapalli:FAA-89-201

[PSA89] Saila Ponnapalli, Tapan K. Sarkar, and Ercument Arvas. Application of conjugate gradient method and FFT to electromagnetics and signal processing problems. In Byrnes and Byrnes [BB90], pages 201–220. ISBN 0-7923-0875-1.

Roy:FAA-89-317

[Roy89] Richard Roy. Invariance techniques in spectral estimation and signal processing. In Byrnes and Byrnes [BB90], pages 317–366. ISBN 0-7923-0875-1.

Sarkar:PSM-91-263

[SA91] T. K. Sarkar and R. S. Adve. Generation of accurate broadband information from narrowband data using the Cauchy method. In Byrnes et al. [BBBH92], pages 263–272. ISBN 0-7923-1804-8.

Saffari:PSM-91-555

[Saf91] B. Saffari. The phase behaviour of ultraflat unimodular polynomials. In Byrnes et al. [BBBH92], pages 555–572. ISBN 0-7923-1804-8.

Schempp:FAA-89-507

[Sch89a] Walter Schempp. Holographic image coding and neurocomputer architectures. In Byrnes and Byrnes [BB90], pages 507–559. ISBN 0-7923-0875-1.

Schroeder:FAA-89-101

[Sch89b] M. R. Schroeder. Number theory and Fourier analysis applications in physics, acoustics and computer science. In Byrnes and Byrnes [BB90], pages 101–130. ISBN 0-7923-0875-1.

Schempp:PSM-91-383

[Sch91] Walter Schempp. Quantum holography and neurocomputer architectures. In Byrnes et al. [BBBH92], pages 383–468. ISBN 0-7923-1804-8.

Seitz:FAA-89-661

[Sei89] K. Seitz. Groupoids and Fourier approximations. In Byrnes and Byrnes [BB90], pages 661–668. ISBN 0-7923-0875-1.

Shapiro:FAA-89-563

[Sha89] H. S. Shapiro. A weighted L^{1} estimate for the Laplace operator. In Byrnes and Byrnes [BB90], pages 563–578. ISBN 0-7923-0875-1.

Suter:PSM-91-217

[SKR^{+}91] Bruce W. Suter, Matthew Kabrisky, Steven K. Rogers, Erik J. Fretheim, Dennis W. Ruck, and Michael Mueller. Linear and nonlinear decomposition of images using Gabor wavelets. In Byrnes et al. [BBBH92], pages 217–233. ISBN 0-7923-1804-8.

Tolimieri:FAA-89-237

[TA89] Richard Tolimieri and Myoung An. Computations in X-ray crystallography. In Byrnes and Byrnes [BB90], pages 237–250. ISBN 0-7923-0875-1.

Tijhuis:FAA-89-273

[Tij89] Anton G. Tijhuis. Fourier methods application in the numerical solution of electromagnetic timedomain scattering problems. In Byrnes and Byrnes [BB90], pages 273–310. ISBN 0-7923-0875-1.

Tuzunalp:FAA-89-261

[Tüz89] Önder Tüzünalp. Auto-cross correlation functions of the even and odd parts of Fourier series. In Byrnes and Byrnes [BB90], pages 261–272. ISBN 0-7923-0875-1.

Unsal:FAA-89-59

[Üns89] Mazhar Ünsal. Complex finite Fourier transform techniques applied to thermal energy storage problems. In Byrnes and Byrnes [BB90], pages 59–72. ISBN 0-7923-0875-1.

Walnut:PSM-91-187

[Wal91] David Walnut. Applications of Gabor and wavelet expansions to the Radon transform. In Byrnes et al. [BBBH92], pages 187–206. ISBN 0-7923-1804-8.